OpenStudy (anonymous):

Will give medal. I need help solving for A where 0

OpenStudy (anonymous):

\[\sin x+\cos x=\sqrt{2}\left( \sin x~\frac{ 1 }{ \sqrt{2} }+\frac{ 1 }{ \sqrt{2}} \cos x~\right)\] \[\cos \frac{ \pi }{ 4 } =\sin \frac{ \pi }{ 4 } =\frac{ 1 }{ \sqrt{2} }\] complete it.